# Robust Data Pruning: Uncovering and Overcoming Implicit Bias

### v1

@@ Appendix A Implementation Details
Our empirical work encompasses three standard computer vision benchmarks (Table 1). All code is implemented in PyTorch [Paszke et al., 2017] and run on an internal cluster equipped with NVIDIA RTX8000 GPUs. We make our code available at https://github.com/avysogorets/fair-data-pruning.
@@
Data pruning methods require different procedures for training the query model and extracting scores for the training data. For EL2N and GraNd, we use $10\%$ of the full training length reported in Table 1 before calculating the importance scores, which is more than the minimum of $10$ epochs recommended by Paul et al. [2021]. To improve the score estimates, we repeat the procedure across $5$ random seeds and average the scores before pruning. Forgetting and Dynamic Uncertainty operate during training, so we execute a full optimization cycle of the query model but only do so once. Likewise, CoreSet is applied once on the fully trained embeddings. We use the greedy k-center variant of CoreSet. Since some of the methods require a hold-out validation set (e.g., MetriQ, CDB-W), we reserve $50\%$ of the test set for this purpose. This split is never used when reporting the final model performance.
@@
We employ data augmentation only when optimizing the final model. The same augmentation strategies are used for all three datasets. In particular, we normalize examples per-channel and randomly apply shifts by at most $4$ pixels in any direction and horizontal flips.
@@
Table 1: Summary of experimental work and hyperparameters. All architectures include batch normalization [Ioffe and Szegedy, 2015] layers followed by ReLU activations. Models are initialized with Kaiming normal [He et al., 2015] and optimized by SGD (momentum $0.9$) with a stepwise LR schedule (0.2×0.2$\times$0.2 × drop factor applied on specified Drop Epochs) and categorical cross-entropy. The above hyperparameters are adopted from prior studies [Frankle et al., 2021, Wang et al., 2020].
@@ B.2 Worst-class optimal priors
We wish to formally establish when $t^{*}(\sigma_{0}/\sigma_{1})$ as defined in Equation 4 minimizes both the average and worst-class risks when $\phi_{0}\propto\sigma_{0}$ and $\phi_{1}\propto\sigma_{1}$. For the first part, as we argued in Section B.1, these priors must satisfy the assumption in Equation 7 (Equation 6 is trivially satisfied). In this case, we can equivalently rewrite it as
@@ B.3 The Effect of SSP
Our setting allows us to adapt and analyze a state-of-the-art baseline pruning algorithm, Self-Supervised Pruning (SSP) [Sorscher et al., 2022]. Recall that, in Section 4, we adopted a variant of SSP that removes samples within a margin $M>0$ of their class means. SSP performs k-means clustering in the embedding space of an ImageNet pre-trained self-supervised model and defines the difficulty of each data point by the cosine distance to its nearest cluster centroid, or prototype. In the case of two 1D Gaussians, this score corresponds to measuring the distance to the closest mean.

### v2

@@ Appendix A Implementation Details
Our empirical work encompasses three standard computer vision benchmarks (Table 1). All code is implemented in PyTorch (Paszke et al., 2017) and run on an internal cluster equipped with NVIDIA RTX8000 GPUs. The total runtime of the empirical work presented in this paper is approximately $45,000$ GPU hours.
@@
Data pruning methods require different procedures for training the query model and extracting scores for the training data. For EL2N and GraNd, we use $10\%$ of the full training length reported in Table 1 before calculating the importance scores, which is more than the minimum of $10$ epochs recommended by Paul et al. (2021). To improve the score estimates, we repeat the procedure across $5$ random seeds and average the scores before pruning. Forgetting and Dynamic Uncertainty operate during training, so we execute a full optimization cycle of the query model but only do so once. Likewise, CoreSet is applied once on the fully trained embeddings. We use the greedy k-center variant of CoreSet. Since some of the methods require a hold-out validation set (e.g., DRoP, CDB-W), we reserve $50\%$ of the test set for this purpose. This split is never used when reporting the final model performance.
@@
We employ data augmentation only when optimizing the final model. The same augmentation strategy is used for all datasets except for Waterbirds where we used none. In particular, we normalize examples per-channel and randomly apply shifts by at most $4$ pixels in any direction and horizontal flips.
Models & Datasets.
As shown in Table 1, we use the following model-dataset pairs: VGG-16 and VGG-19 (Simonyan & Zisserman, 2015) on CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), respectively, ResNet-18 (He et al., 2016) on TinyImageNet (MIT License) (Le & Yang, 2015), ImageNet pre-trained ResNet-50 on Waterbirds (Sagawa* et al., 2020) (MIT License), and ResNet-50 on ImageNet (Deng et al., 2009).
@@
ResNet-50 ImageNet $90$ $30/60$ $512$ $0.2$ $1e$-$4$
ResNet-50 Waterbirds $300$ None $128$ $0.0001$ $1e$-$4$
Table 1: Summary of experimental work and hyperparameters. All architectures include batch normalization (Ioffe & Szegedy, 2015) layers followed by ReLU activations. Models are initialized with Kaiming normal (He et al., 2015) and optimized by SGD (momentum $0.9$) with a stepwise LR schedule (0.2×0.2$\times$0.2 × drop factor applied on specified Drop Epochs) and categorical cross-entropy. These hyperparameters are adopted from prior studies (Frankle et al., 2021; Wang et al., 2020; Sagawa* et al., 2020; Cai et al., 2022).
@@ B.2 Worst-class optimal priors
We wish to formally establish that $t^{*}(\sigma_{0}/\sigma_{1})$ as defined in Equation 4 minimizes both the average and worst-class risks when $\phi_{0}\propto\sigma_{0}$ and $\phi_{1}\propto\sigma_{1}$. For the first part, as we argued in Section B.1, these priors must satisfy the assumption in Equation 7 (Equation 6 is trivially satisfied). In this case, we can equivalently rewrite it as
@@ B.3 The Effect of SSP
Our setting allows us to adapt and analyze a state-of-the-art baseline pruning algorithm, Self-Supervised Pruning (SSP) (Sorscher et al., 2022). Recall that, in Section 3, we adopted a variant of SSP that removes samples within a margin $M>0$ of their class means. SSP performs k-means clustering in the embedding space of an ImageNet pre-trained self-supervised model and defines the difficulty of each data point by the cosine distance to its nearest cluster centroid, or prototype. In the case of two univariate Gaussians, this score corresponds to measuring the distance to the closest mean.
@@ B.4 Multivariate Isotropic Gaussians
Appendix C Data Pruning: Full Results
Figure 10: The average test performance of various data pruning algorithms against dataset density (fraction of samples remaining after pruning) and metrics of class robustness. Top to Bottom: VGG-16 on CIFAR-10; VGG-19 on CIFAR-100; ResNet-18 on TinyImageNet; ResNet-50 on ImageNet. All results averaged over $3$ random seeds. Error bands represent min/max.
Appendix D DRoP: Full Results
(a) VGG-16 on CIFAR-10
(b) VGG-19 on CIFAR-100
(c) ResNet-18 on TinyImageNet
Figure 11: The average test performance of various data pruning protocols against dataset density and measures of class robustness. Random+DroP consistently outperforms all baselines with respect to these additional measures of distributional robustness besides worst-class accuracy reported in Figure 5, too, confirming its effectiveness in reducing bias. All results averaged over $3$ random seeds. Error bands represent min/max.
Figure 12: The average test performance of various data pruning protocols against dataset density and measures of class robustness. ResNet-50 on ImageNet with EL2N as Strategy. All results averaged over $3$ random seeds. Error bands represent min/max.
Appendix E Data Pruning: Class-Wise Analysis
Appendix F DRoP Ablation: Pre-Training Length
As mentioned in Section 4, DRoP requires a pre-trained query model to compute class-wise validation errors. All existing data pruning methods also require pre-training to compute sample-wise pruning scores. However, as Paul et al. (2021) shows, it suffices to run optimization for as little as $10$ epochs to achieve accurate enough scores for EL2N. In our experimental work, we used $10\%$ of the full training cycle to train the query model for EL2N, GraNd, and DRoP, which is at least the recommended $10$ epochs for all model-dataset pairs tested in this study. In this section, we verify that DRoP reaches its full performance with this much pre-training and, hence, it does not introduce any additional computational overhead already needed by the current data pruning algorithms. Indeed, Figure 15 shows that the downstream performance of DRoP saturates at $10$ pre-training epochs or earlier, with lower pruning ratios allowing even less training than more aggressive ones.
Figure 15: Final model performance after DRoP with different lengths of the query model training. Top: VGG-19 on CIFAR-100, Bottom: ResNet-18 on TinyImageNet. Repeated over $3$ random seeds and three dataset densities ($d=0.4,0.5,0.8$). In our study, we use $16$ and $9$ pre-training epochs for VGG-19 on CIFAR-100 and ResNet-18 on TinyImageNet (10% of the full training cycle), respectively.